zombieblud:

There is a daily fee for renting a moving truck, plus a charge of $0.50 per mile driven. If driven 48 miles, it costs $64 to rent a truck. Write and solve and equation to find the daily fee.

Luigi0210:

So we have the equation \( ax+b =c \) where a - charge per mile x - miles driven b - daily fee c - total cost Plugging everything in you get: \(0.50x + b = 64 \) It is given that they drove 48 miles, so you can plug that in as well and solve: \(0.50(48)+b = 64 \) \( 24+b = 64\) And you get b as \( b= 40 \) So the daily fee for renting the moving truck is \(\color{gold}{$40}\) (Assuming this is just a one-day use).

zombieblud:

@luigi0210 wrote:
So we have the equation \( ax+b =c \) where a - charge per mile x - miles driven b - daily fee c - total cost Plugging everything in you get: \(0.50x + b = 64 \) It is given that they drove 48 miles, so you can plug that in as well and solve: \(0.50(48)+b = 64 \) \( 24+b = 64\) And you get b as \( b= 40 \) So the daily fee for renting the moving truck is \(\color{gold}{$40}\) (Assuming this is just a one-day use).
Thank you!!!